# SoK: Payment Channel Networks

### v2

@@ VI Rebalancing
Motivation:Rebalancing in payment channel networks involves redistributing the balance within an existing payment channel or across multiple channels to maintain or improve the network’s liquidity and usability. This is done to ensure that both parties in the channel have sufficient funds on their respective sides to send and receive payments. Rebalancing is crucial to payment channel networks. If the link weights of nodes become zero/depleted as a result of being involved in transactions, it will prevent them from taking part in further payments until those link weights are replenished. Such nodes are called as dormant nodes in the network. A PCN which has a large number of such dormant nodes experiences an overall reduction in transaction throughput and an overall reduction in liquidity. Hence, the link weights of the nodes in the PCN need to be rebalanced to prevent failure of transactions due to lack of liquidity. A rebalancing protocol should be be able to replenish the link weight of dormant nodes in the PCN irrespective of the PCNs topology, while not having to employ any trusted entity and should not compromise privacy of nodes. We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of rebalancing protocols in Table V. We describe the metrics as follows:
Ideal properties:
1) Blockchain: This is the blockchain on which the corresponding rebalancing protocol can be deployed. As it can be inferred from the Table, the protocols can either be deployed on the Ethereum or the Bitcoin blockchains.
2) Trusted entity: This metric determines the nature of trust required among parties to deploy the rebalancing protocol. A centralized trust (C) indicates the presence of a single trusted entity and distributed trust (D) indicates the distribution of trust across various entities.
3) Graph compatibility: This indicates the nature of the topology of the network on which a rebalancing protocol can be deployed. Cycles indicate that the rebalancing protocol can only deployed on cyclic graph and agnostic indicates that the rebalancing protocol can be deployed irrespective of the network topology.
4) Privacy: Privacy is said to be achieved when sensitive information such as the local balance and the identities of nodes are not known to anyone except for their immediate neighbors.
@@ VII Research Gaps & Open Problems
In this section, we highlight the gaps in research published until now in the areas of pathfinding and routing, virtual channel construction, payment channel hubs, and state channels. This section provides an overview of the current research problems in these areas along with a discussion on why is developing a solution to these problems an interesting research challenge.
VII-A Gaps in pathfinding and routing for PCNs
We now identify the gaps in research in pathfinding and routing protocols for payment channel networks.
@@ VII-A Gaps in pathfinding and routing for PCNs
1) Super node liquidity validation in PCNs: A super node is a highly connected node (it has payment channels open with numerous other nodes in the network) and also has a high liquidity in those channels. These super nodes aid other nodes in pathfinding and routing of payments. These super nodes are called as a trampoline nodes in $\mathsf{LN}$ [89, 90, 91, 92]. Several other works in the literature refer to super nodes as routing node, routing helper, landmark node, routers, etc [25, 24, 27, 37, 33]. One of the main problems with the super nodes is that the sender has no way to know whether the super node possess the local balance (liquidity) to route the payment of the sender. The local balance of a super node or any node in a PCN is a private value and should not be known to any node except for its immediate neighbor. Currently, if a super node does not have enough liquidity to route the payment of a sender, the transaction fails and it has to be retried by the sender. In $\mathsf{LN}$, one of the most widely used PCN, this is a significant problem. Sometimes the sender might have to keep retrying for $\approx$ 1 hour to have a successful payment [93]. The main goal of PCNs is to facilitate instantaneous payments and these transaction retries render such payments almost impossible. It might greatly benefit the sender if the it has a mechanism to validate whether a super node has enough liquidity (balance) to route its payment without violating any privacy concerns in the network. Achieving this is hard because the local balance a node has with its immediate neighbors in a PCN is considered private information and the sender has to know whether the super node possess enough liquidity without comprising privacy.
@@
2) Channel verification in a PCN: To be a part of any PCN, nodes will open payment channels with other nodes in the network. Two nodes open a payment channel between them by posting a transaction to the blockchain. This transaction can be posted on the blockchain (as in UTXO based blockchains like BTC) or this transaction can be posted as a function call to an existing smart contract (as in blockchains that support smart contracts like ETH). In the most popular PCN, $\mathsf{LN}$, the channel verification procedure is very inefficient. It involves the node who intends to verify a channel to request the block in which the channel opening transaction has been included, verifying whether the transaction has been successfully executed by the validator/miner and finally verifying if the channel opening transaction corresponds to a 2-2 multi signature address on the blockchain. The verifier performing these steps is inefficient since all these steps will have already been performed by the miner. According to the current verification process, any node would need to perform all of these steps for each payment channel it has open in the network. Consider the following scenario: a node $j$ in the PCN has payment channels open with $m$ nodes in the PCN. Furthermore, assume that each neighbor opens $k$ new channels for each time epoch. For this example, assume the value of the time epoch is 1 block. This would mean that the total number of channels the node $j$ has to verify a total of $m*k$ channels for every block. This requires the node $j$ to access the blockchain a total of $m*k$ times to perform channel verification. Finding a way to do this without blockchain access and in a blockchain agnostic manner in a hard research challenge.
VII-B Gaps in Tumblers
We will now describe the research gaps in tumblers. 1) Privacy against aborts: In a given a transaction, the tumbler should not be able to correctly identify the sender/receiver pair involved. Currently, there is only one construction [59] that guarantees this. However, this construction works only for blockchains that support Turing complete programming languages. Currently, there is no such construction for UTXO based blockchains that guarantees privacy against aborts. This is hard because if the sender aborts a transaction, the tumbler will know which receiver did not receive the payment. This is hard because, a tumbler mixes multiple transactions together to ensure unlikability among payments. Also, a tumbler should enforce atomicity so that the payment either goes through completely or does not go through at all. In the case of an abort, an adversary who controls the tumbler might be able to deduce which receiver amongst the ones involved did not receive a payment and will be able to link the corresponding transaction amount to that receiver.
2) Time epochs: A tumbler should be able to facilitate payments without having a predefined time interval or epoch. Currently, the majority of payment channel hub constructions process payments in discrete epochs to increase the size of their anonymity sets. An anonymity set refers to a set of users (senders/receivers) whose identity is indistinguishable from each other. The size of this set is proportional to the privacy of sender/receiver in the transaction. When the interval between two consecutive time epochs is sufficiently large, it will promote more users to take part in the transactions within a time epoch, which improves privacy. Not having time epochs results in smaller anonymity sets and hence could possibly weaken privacy. The trade off is between privacy and flexibility.
3) Dynamic corruption: A tumbler should ideally be resistant to dynamic corruption. This means the protocol should be resilient to parties deviating from their ideal execution during the beginning, progress, and ending of the protocol’s execution. Current tumbler constructions only provide support for static corruption, which may be unrealistic from a practical standpoint. Security against dynamic corruption is hard to achieve because the tumbler mixes several transactions together to preserve anonymity. Since the tumbler should not be able to link transaction amount to a sender/receiver pair (ideally), if the sender/receiver becomes malicious during the protocol’s execution, ideally, that transaction must be aborted to prevent financial loss to the honest party. However, the tumbler should not be able to link a transaction to a sender/receiver pair, making security against dynamic corruption a hard research problem.
VII-C Gaps in Virtual Channels
We will now describe the open problems in the area of virtual channels.
1) Blockchain agnostic: A virtual channel should ideally be blockchain agnostic, meaning it should be deployable regardless of the underlying scripting or programming language structure supported by the blockchain.
@@ VII-C Gaps in Virtual Channels
2) Well defined fee structure: The intermediary/intermediaries involved in the virtual channel construction additionally lock coins apart from the ones locked in the underlying payment channel. It is necessary to have a well-defined fee structure for this additional service. Currently, nodes get paid a routing fees/forwarding fees for every transaction they process. This fee is dependent on the transaction amount being processed. In the case of virtual channels, having a well-defined fee structure is difficult due to the following reasons: 1) The fee structure should take into account the amount of funds and the time for which these funds of the intermediary/intermediaries are locked in virtual channel. 2) It also needs to take into account the routing/forwarding fees an intermediary could have earned by not locking up the coins in the virtual channel. The fees for an intermediary should be ideally be more than the routing fees it would have earned in order to motivate nodes to be involved in virtual channel construction. Currently, no virtual channel construction has a well-defined fee structure.
3) Off-chain dispute resolution: Ideally a virtual channel protocol should provide the flexibility to resolve any disputes without accessing the blockchain. We have seen in Section I, that blockchain writes are expensive. Having an off-chain dispute resolution is a hard research challenge due to the following reasons; 1) Any form of layer-2 disputes are usually resolved on the blockchain. Since underlying consensus in blockchain usually guarantees resolution of a dispute raised within a specific time period. In an off-chain scenario, there is no such mechanism which guarantees this. None of the virtual channel constructions have a provision of off-chain dispute resolution.
4) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by constructing a single virtual channel from the sender to the receiver, without establishing virtual channel(s) between any pair of intermediate nodes along the path. Currently, no VC construction has this capability. This is hard because, multihop virtual channel involves designing a mechanism for sequentially locking coins across a path comprising of several intermediate nodes. Furthermore, the construction of a multihop virtual channel should also ensure that any transaction does not effect the overall balance neutrality (no node apart from the receiver along the path gains or loses any money). Also, the goals of privacy, atomicity, offchain dispute resolution need to be enforced.
VII-D Gaps in state channel protocols
We will now describe the gaps in the area of state channel protocols.
1) General purpose state channels: An ideal blockchain should be able to facilitate deployment and execution of several decentralized applications without requiring the users to access the blockchain for any purpose other than opening and closing the state channel. Currently, there are two such constructions [49, 48] which incorporate code reusability into their design. These two constructions support UTXO based blockchains and have no provision for offchain dispute resolution.
@@ VII-D Gaps in state channel protocols
2) Offchain dispute resolution: Current state channel protocols use an onchain transaction or a function call to the onchain smart contract in the case of dispute resolution. An ideal state channel protocol should be able to facilitate dispute resolutions in an offchain manner. This is a hard research challenge since disputes on the blockchain are usually resolved by the underlying consensus mechanisms. In an offchain scenario, there is no such consensus that guarantees transaction validity. Also, in an offchain protocol such as the state channel, nodes which are not part of the state channel protocol do not know any details of the protocol’s execution due to privacy concerns. Onchain transactions on the hand are publicly accessible.
VIII Network Statistics
(a) Minimum accepted value of HTLC payment (msat)
(b) Distribution of base fee (msat)
(c) Distribution of rate fee (msat)
(d) Distribution of node degree
(e) Distribution of HTLC expiration time
Figure 5: $\mathsf{LN}$ statistics
This section introduces a general analysis of the parameters used in $\mathsf{LN}$, extracted from $\mathsf{LN}$ gossip messages [94]. It offers views into user behavior and the implications of such behavior for network performance. Our analysis shows that the nodes prefer to choose the default value of the $\mathsf{LN}$ parameters. Using standard default values for these parameters allows nodes to operate smoothly together, reducing the chance of compatibility issues and enhancing the overall reliability and efficiency of the $\mathsf{LN}$. We evaluated the analysis on a snapshot of $\mathsf{LN}$ topology from November 27, 2021 [95, 96]. The snapshot includes 18,331 nodes and 80,918 channels with a mean channel capacity of 4056152.35 satoshi and a median channel capacity is 1000000 satoshi (this provides us information on how funds are distributed among $\mathsf{LN}$ channels).
Although not explicitly shown in this chart, nodes with multiple connections suggest the presence of fewer central nodes compared to the end-users nodes. Figure 5(a) shows that the majority of the nodes (87.9%) set the ’htlc_minimum_msat’ to a default value of 1000 msat, which indicates that the $\mathsf{LN}$ supports micropayment transactions. Figure 5(b) shows that ’fee_base_msat’ parameter across all of the nodes has the highest two values 1000 msat as a default value, and 0 msat as the minimum value, which indicates that most of the nodes do not set the fee_base_msat beyond the default configuration, and likely to support the network usage. Figure 5(c) shows the majority of the nodes (51.8%) in $\mathsf{LN}$ use various ’fee_rate_milli_msat’ values. A substantial portion of nodes (35.2%) set their fee rate to 1 msat to attract more transactions. Also, different nodes have different fee rate strategies, making $\mathsf{LN}$ a competitive environment. Figure 5(d) shows that the majority of the nodes (49.5%) have only one connection that indicates that in $\mathsf{LN}$, many end-users connected to fewer well-connected nodes. Moreover, 30% of the nodes have no payment channels, these could be new nodes added to the $\mathsf{LN}$ or they could be inactive nodes. Figure 5(e) shows that (81.5%) of nodes prefer to set the ’cltv_expiry_delta’ to the default value of 40 blocks, which is the number of the blocks a node can wait before losing its satoshis from $\mathsf{LN}$. Moreover, it helps to increase the flexibility of the $\mathsf{LN}$ to include more nodes along the payment path before reaching the ”locktime max’ parameter value (2016 blocks). According to the analysis, the nodes generally prefer to use the default parameter values of the lightning network. Using these default values can help in developing more effective routing protocols and optimizing routing fee mechanisms for the nodes.
IX Conclusion
@@
In this paper, we qualitatively compared the recent work in various areas of PCN research: pathfinding and routing, virtual channels, payment channel hubs, state channels and rebalancing. We have also provided the gaps in research in these areas along with reasons why fulfilling those gaps is a hard research challenge. We hope that this paper motivates researchers to build robust protocols that address these gaps that would go a long way towards building out and developing the decentralized finance ecosystem.

### v3

@@ VI Rebalancing
Motivation: Rebalancing in PCNs involves redistributing the balance within an existing payment channel or across multiple channels to maintain or improve the network’s liquidity and usability. This is done to ensure that both parties in the channel have sufficient funds on their respective sides to send and receive payments. Rebalancing is crucial to payment channel networks. If the link weights of nodes become zero/depleted as a result of being involved in transactions, it will prevent them from taking part in further payments until those link weights are replenished. Such nodes are called as dormant nodes in the network. A PCN which has a large number of such dormant nodes experiences an overall reduction in transaction throughput and an overall reduction in liquidity. Hence, the link weights of the nodes in the PCN need to be rebalanced to prevent failure of transactions due to lack of liquidity. A rebalancing protocol should be be able to replenish the link weight of dormant nodes in the PCN irrespective of the PCNs topology, while not having to employ any trusted entity and should not compromise privacy of nodes. We now give the properties desired from an ideal rebalancing protocol, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of rebalancing protocols in Table V. We describe the metrics as follows:
Ideal properties: 1) Trusted entity: This metric determines the nature of trust required among parties to deploy the rebalancing protocol. Centralized trust (C) indicates the presence of a single trusted entity and distributed trust (D) indicates the distribution of trust across various entities. Importance: It is ideal to have distributed trust for a rebalancing protocol since, PCN payments are by nature decentralized. 2) Graph compatibility: This indicates the nature of the topology of the network on which a rebalancing protocol can be deployed. Cycles indicate that the rebalancing protocol can only deployed on cyclic graphs and agnostic indicates that the rebalancing protocol can be deployed irrespective of the network topology. Importance: Ideally a rebalancing protocol should be deployable irrespective of the topology of the graph since its practically infeasible to assume that every node in a PCN will be a part of a cycle. 3) Privacy: Privacy is said to be achieved when sensitive information such as the local balance and the identities of nodes are not known to anyone except for their immediate neighbors. Importance: The importance of privacy for layer-2 protocols has already been described in Section II.
Rebalancing is a relatively mature area and significant challenges have already been addressed.
@@ VII Research Gaps & Open Problems
In this section, we highlight the gaps in research published up until now in the areas of pathfinding and routing, virtual channel construction, state channels and payment channel hubs. The gaps are described as research questions, denoted by RQ.
@@
RQ1: Why is super node liquidity validation in PCNs hard? A super node, variously called as a trampoline node, routing node, routing helper, landmark node, router, etc. [48, 45, 54, 64, 62] is a highly connected node with numerous high liquidity channels, that helps in pathfinding and routing payments. One of the main problems with the super nodes is that a sender has no way to know whether the super node possesses enough liquidity on its channels to route a payment. The local balance of a super node in a given channel (or of any node in a PCN) is a private value and should not be known to any node except for its immediate neighbor that it shares the channel with. Currently, if a super node does not have enough liquidity to route the payment of a sender, the transaction fails and it has to be retried by the sender. In $\mathsf{LN}$, one of the most widely used PCN, this is a significant problem. Sometimes the sender might have to keep retrying for $\approx$ 1 hour to have a successful payment [111]. The main goal of PCNs is to facilitate instantaneous payments and these transaction retries render such payments almost impossible. It will greatly benefit the sender if it has a mechanism to validate whether a super node has enough liquidity (balance) to route its payment without violating any privacy concerns.
@@
RQ2: Why is channel verification in a PCN hard? To be a part of any PCN, nodes will open payment channels with other nodes in the network. Two nodes open a payment channel between them by posting a transaction to the blockchain. This transaction can be posted on the blockchain or as a function call to an existing smart contract. In the most popular PCN, $\mathsf{LN}$, the procedure of verifying whether a payment channel really exists on the blockchain is very inefficient. A node who wants to verify a channel needs to request the block in which the channel opening transaction has been included, verifying whether the transaction has been successfully executed by the validator/miner and finally verifying if the channel opening transaction corresponds to a 2-2 multi signature address on the blockchain. The verifier performing these steps is inefficient since all these steps will have already been performed by the miner. Finding a way to do this without blockchain access and in a blockchain agnostic manner in a hard research challenge.
RQ3: Why is designing pathfinding protocols for PCNs, that comprise of several distinct well-connected components a hard problem? Though solutions such as [62] exist that solve this problem to a certain extent by using routing helpers/trampoline nodes, the aspect that makes it hard is to quantify the denseness/sparseness of a well-connected component. If we assume that we deploy one trampoline node for each well-connected component, there may be well-connected components in the PCN which only have a few nodes (i.e., islands). Selection of a trampoline node amongst these nodes is difficult.
@@
RQ4: Why is designing a routing protocol that supports concurrent payments and is resilient to channel gaming a hard problem? Processing concurrent transactions requires the design of a mechanism that allows a node to lock a portion of its liquidity in a channel with an immediate neighbor for one transaction while simultaneously using the remaining liquidity to process another. Though there are protocols that support concurrency [47, 45, 48, 62, 63, 64], they are not resilient to the presence of potentially malicious nodes in the PCN, which may initiate transactions with the sole intent of locking liquidity, leading to congestion and disruption in the network. RQ5: Why is having a well-defined fee structure for virtual channels hard? The intermediary(s) involved in the virtual channel construction additionally lock coins in virtual channels apart from the ones locked in the underlying payment channel. Currently, nodes get paid routing fees for every transaction they process. In the case of virtual channels, having a well-defined fee structure is difficult due to the following reasons: 1) The fee structure should take into account the amount of funds and the time for which these funds of the intermediary(ies) are locked in a virtual channel. 2) It also needs to take into account the routing fee an intermediary could have earned by not locking up the coins in the virtual channel.
RQ6: Why is off-chain dispute resolution in virtual channels hard? There is no offchain consensus mechanism for dispute resolution in a PCN, as opposed to the 51% honest majority assumption that exists among validators on the blockchain. This honest majority helps resolve disputes in the transactions posted to the blockchain. Designing such a dispute mechanism for layer-2 transactions is hard since transactions are private (not posted to the blockchain), and nodes do not broadcast their activities to the entire network.
RQ7: Why is providing support for a multihop virtual channel a hard problem? This is hard since a multihop virtual channel construction should ensure that neither sender/receiver nor the intermediate nodes should lock coins in multiple channels at the same time.
RQ8: Why is ensuring privacy in a virtual channel protocol a hard problem? In a recursive virtual channel, new virtual channels are constructed on top of existing virtual channels to facilitate payments. This staggered nature makes it mandatory to reveal the identity of at least one endpoint node (sender or receiver). This is because, at least one node among the sender/receiver is involved in all virtual channels. The solution to this problem is to design a multihop virtual channel.
RQ9: Why is designing a payment channel hub (PCH) that is resistant to privacy against aborts and dynamic corruption a hard challenge? PCHs usually use transaction mixing for enhancing privacy, which is a process in which multiple payments from different users are mixed together in such a way that it is infeasible for the hub to link the sender and recipient of a specific transaction. This process helps obscure the flow of funds, providing unlinkability. Designing a payment channel hub that is resistant to privacy against aborts is hard because, if a PCH selectively aborts a payment from a sender/receiver, the counter party whose payment also failed can be linked. If sender/receiver gets corrupted during the PCH’s execution, the corresponding transaction has to be aborted to ensure atomicity, which is why the existing tumbler constructions assume a static adversary, in which certain nodes are designated as corrupted before the PCH begins execution. The trade off here is preserving transaction unlikability during a corrupted party’s transaction abort.
@@
RQ10: Why is having an offchain dispute resolution for a state channel hard? Current state channel protocols use an onchain transaction or a function call to the onchain smart contract in the case of dispute resolution. An ideal state channel protocol should be able to facilitate dispute resolutions in an offchain manner. This is a hard research challenge since disputes on the blockchain are usually resolved by the underlying consensus mechanisms. In an offchain scenario, there is no such consensus that guarantees transaction validity. Also, in an offchain protocol such as the state channel, nodes which are not part of the state channel protocol do not know any details of the protocol’s execution due to privacy concerns. Onchain transactions on the hand are publicly accessible.
VIII Conclusion
@@
In this paper, we qualitatively compared the recent work in various foundational areas of PCN research: pathfinding and routing, virtual channels, state channels, payment channel hubs, and rebalancing protocols. We also discussed the gaps in research in these areas along with reasons why fulfilling those gaps is non-trivial. We hope that this paper motivates researchers to build robust protocols that address these gaps that would go a long way towards building out and developing a decentralized financial ecosystem.
Acknowledgement
This material is based upon work supported by the National Science Foundation under Award No. 2148358, 1914635, 2417062, and the Department of Energy under Award No. DE-SC0023392. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation and the Department of Energy.